OpenStudy (anonymous):

Find the probability distribution function for the number of jazz CDs when 4 CDs are selected from a collection consisting of 5 Jazz CDs, 2 classical CDs, and 3 rock CDs. Express your results by means of a formula.

OpenStudy (anonymous):

let x = number of jazz cd's selected. then x = {0, 1, 2, 3, 4} \[P(X=k)=\frac{ \left(\begin{matrix}5 \\ k\end{matrix}\right)\left(\begin{matrix}5 \\ 4-k\end{matrix}\right) }{ \left(\begin{matrix}10 \\ 4\end{matrix}\right) }\]